# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v2

@@ 4 Results
Comparison with PEFT CL approaches: As PEARL entails dynamic allocation of LoRA components, we compare PEARL with state-of-the-art PEFT CL methods across 5, 10, and 20 tasks in ImageNet-R [27], in Table 2. The experiment is conducted using the ViT-B/16 backbone [16] pre-trained on ImageNet-21k. As can be seen, PEARL (ViT) shows a significant improvement in both incremental and final accuracy, with absolute gains ranging from 9% to 14%. The performance improvement is more pronounced in 5 task sequence and slightly decreases thereafter as we move towards 10 and then to 20 task sequences. Among considered baselines, InfoLoRA is the closest work in terms of design and performance, but falls short of PEARL (ViT)’s performance by a huge margin across all scenarios. InfLoRA injects a small number of parameters to re-parameterize the pre-trained weights and shows that finetuning these injected parameters is equivalent to finetuning the pre-trained weights within a subspace. Since the reduction matrix $B$ infLoRA is handcrafted rather than being learnt from the training data, InfLoRA suffers a performance degradation compared to PEARL (ViT). Additional comparisons between our method and adapter-based as well as prompt-based approaches can be found in Table 14 and Table 14.
@@ A.2 Justification for considered baselines
To facilitate extensive evaluation of PEARL, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE); (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM); (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE); and (iv) PEFT approaches for continual learning (e.g., L2P, C-LoRA, InfLoRA). Firstly, inspired by the GWT, exemplar-based baselines entail a shared network for information reuse and task-specific attention modules that re-project the representation from the common representation space to the global workspace. Likewise, the PEARL framework includes a shared reference task network for information reuse and task-specific, layer-specific low-rank adapters for orienting towards the current task. Secondly, PEARL is a rehearsal-free framework; it does not store previous task samples either directly or indirectly through generative networks. Therefore, it aligns well with the NECIL benchmark and associated approaches. Thirdly, we note that low-rank adaptation introduces task-specific parameters that are trained on the corresponding task and kept frozen afterwards, akin to parameter-isolation approaches. As PEARL exhibits partial parameter isolation with a shared reference task network and low-rank adapters, we also provide a comparison with parameter isolation approaches. Finally, since PEARL readily utilizes LoRA [29], a prominent PEFT technique, we offer an in-depth comparison with PEFT continual learning approaches. In addition, we also provide a discussion on related dynamic low-rank adaptation techniques beyond continual learning in Section A.3.
@@ A.3 Discussion on PEFT LoRA counterparts
PEFT CL approaches with static rank allocation: Among PEFT continual learning approaches, InfLoRA [45] and C-LoRA [59] are prominent techniques that involve low-rank adaptation of pre-trained networks, similar to PEARL. Each of these methods are tailored for CL scenarios and addresses unique challenges. InfLoRA aims to minimize the interference of new tasks on previously learned ones, thereby improving knowledge retention. However, its reliance on a handcrafted reduction matrix $B$, rather than a data-driven learned representation, limits its flexibility. C-LoRA, on the other hand, depends on a pre-trained model as its foundation, making it reliant on the availability of such models for its application. It applies LoRA to both the key and query attention mechanisms, ensuring efficient adaptation for new tasks. Additionally, C-LoRA incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. Despite its relative success, caution is warranted when training C-LoRA on large task sequences, as this approach suffers from scalability issues. In line with standard practices in continual learning benchmarks, we provide results for SeqLoRA, a sequential finetuning baseline without any safeguards against catastrophic forgetting. SeqLoRA experiences significant forgetting, underscoring the difficulty of maintaining performance across tasks in continual learning.
@@
PEFT approaches with dynamic rank allocation: Parameter efficient fine-tuning techniques such as LoRA distribute trainable parameters equally across all layers, which may reduce efficiency when layers vary in their contribution towards task adaptation. To address this, adaptive low-rank adaptation strategies have emerged that dynamically adjust parameter budgets to improve fine-tuning efficiency, similar to PEARL. AdaLoRA [76] adaptively allocates the parameter budget among weight matrices according to their importance score. The importance score is computed by parameterizing the incremental updates in the form of SV-decomposition and the singular values of unimportant updates are pruned. DyLoRA [63] proposed a dynamic low-rank adaptation in which LoRA blocks are trained for a range of ranks instead of a single rank by sorting the representation learned by the adapter module at different ranks during training. More recent methods, such as DoRA [50] and Dynamic Rank-Selective LoRA [48], have emerged, with DoRA decomposing high-rank LoRA layers into structured single-rank components, while Dynamic Rank-Selective LoRA adaptively assigns ranks to LoRA modules based on their relevance to the current data in CL in vision language models. Although these approaches address a similar challenge as PEARL, these are not tailored for CL in custom, compact models.
@@ B.1 Comparison with fixed rank as a baseline
The PEFT CL approaches offer lightweight training schedules while effectively retaining the consolidated knowledge of pre-trained models. However, LoRA components in these methods are particularly sensitive to rank selection, which can lead to suboptimal resource allocation and diminished performance [62]. Table 6 summarizes the average model size and performance metrics across various ranks for LoRA components. As the rank increases, performance improves at the expense of increased model size; however, beyond a certain rank, the performance gains do not really justify the additional increase in model size. Consequently, tuning the rank is essential to strike the right balance in this trade-off where model growth remains modest with optimum performance. To this end, PEARL employs a task-specific, layer-specific dynamic rank allocation enabling an optimal trade-off between model size and performance across different datasets and architectures.
@@ B.2 Comparison with convolutional parameter isolation approaches
Figure 2: Task-IL performance comparison with parameter isolation approaches on Seq-CIFAR100 (20T). We report the final Task-IL accuracies of each task after training on all tasks. Methods compared are NISPA [26], PNN [57], PAE [31], CPG [30], PackNet [49], and CLNP [23]
@@ B.5 Hyperparameter Tuning
Table 10: A benchmark comparison with CNN-based prior works on Class-IL ($A_{T}$) and Task-IL ($A_{T}$). The best results are highlighted in bold while the second best are underlined. The error terms indicate a standard deviation across 3 random runs. The rehearsal-based baselines employ a buffer size of 200.
@@
Class-IL ($A_{T}$) Task-IL ($A_{T}$) Class-IL ($A_{T}$) Task-IL ($A_{T}$)
@@
PNNs [57] NeurIPS 2016 - 74.01 $\pm$1.11 - 67.84 $\pm$0.29
PackNet [49] CVPR 2018 - 72.39 $\pm$0.37 - 60.46 $\pm$1.22
NISPA [26] NeurIPS 2022 - 65.36 $\pm$2.19 - 59.56 $\pm$0.32
SparCL-EWC [69] NeurIPS 2022 - 59.53 $\pm$0.25 - 59.56 $\pm$0.32
@@
DER++ [10] NeurIPS 2020 29.60 $\pm$1.14 62.49 $\pm$1.02 10.96 $\pm$1.17 40.87 $\pm$1.16
ER-ACE [11] ICLR 2022 35.17 $\pm$1.17 63.09 $\pm$1.23 11.25 $\pm$0.54 44.17$\pm$1.02
Co2L [12] ICCV 2021 31.90 $\pm$0.38 55.02 $\pm$0.36 13.88 $\pm$0.40 42.37 $\pm$0.74
CLS-ER [2] ICLR 2022 43.80 $\pm$1.89 73.49 $\pm$1.04 23.47 $\pm$0.80 49.60 $\pm$0.72
OCDNet [42] IJCAI 2022 44.29 $\pm$0.49 73.53 $\pm$0.24 17.60 $\pm$0.97 56.19 $\pm$1.31
TAMiL [7] ICLR 2022 41.43 $\pm$0.75 71.39 $\pm$0.17 20.46 $\pm$0.40 55.44 $\pm$0.52
IMEX-Reg [4] TMLR 2024 48.54 $\pm$0.23 75.61 $\pm$0.73 24.15 $\pm$0.78 62.91 $\pm$0.54
TriRE [64] NeurIPS 2023 43.91 $\pm$0.18 71.66 $\pm$ 0.44 20.14 $\pm$0.19 55.95 $\pm$0.78
AGILE [3] CoLLAs 2024 45.73 $\pm$0.15 74.37 $\pm$0.34 20.19 $\pm$1.65 53.47 $\pm$1.60
SARL [58] ICLR 2025 48.96 $\pm$ 0.53 78.91 $\pm$ 0.15 28.95 $\pm$ 1.13 71.56 $\pm$0.18
@@ B.7 Performance in longer task sequences
Figure 3: Evaluation of Class-IL ($A_{T}$) performance under longer task sequences in Seq-CIFAR100. In line with other parameter isolation approaches, there is a steady drop in performance as we move from 5 tasks to 10 tasks and eventually to 20 tasks.
@@ B.9 Dynamic rank allocation across layers in ViT
PEARL introduces a PEFT CL framework that incorporates dynamic rank selection for LoRA components during training. PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 1 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T). From left to right, Figure 1 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Similarly, Figure 4 presents the dynamic rank allocation across layers for the final task in ImageNet-R (5T) for PEARL (ViT). In the ViT model, earlier layers are primarily responsible for extracting low-level image features. These features lay the foundation for subsequent high-level feature extraction and semantic understanding in later layers. As these features relatively less complex and can be re-used across tasks, PEARL puts less emphasis on them through lower rank allocation. On the other hand, intermediate layers handle more abstract and complex features, such as parts of objects and entire objects, necessitating higher rank allocation. Therefore, PEARL puts a major emphasis on intermediate layers with higher rank allocation. The top layers of the ViT model are designed to capture high-level semantic features and global contextual information. By later layers, the model has likely already captured the most significant semantic features relevant to the task, and additional layers contribute marginally smaller improvements to the final output quality. Therefore, PEARL allocates a lower threshold and a corresponding lower rank in later layers. In line with the aforementioned findings in Zeng et al. [75], PEARL displays a bell-curve shaped dynamic rank allocation across layers thereby optimizing resource allocation and maximizing information reuse.
@@ B.13 Additional comparison with prompt-based methods
Table 15 provides an extensive comparison of prompt-based methods in the continual learning (CL) setting on ImageNet-R. Following HiDe-Prompt [66], we evaluate our method using the final forgetting measure (FFM) and cumulative forgetting measure (CFM), defined as follows. Let $A_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task.
@@
Table 15: Following [66], we compare additional prompt-based methods in the Class-IL setting on ImageNet-R using a ViT-B/16 backbone pre-trained on ImageNet-21K. Bold indicates the best result and underline indicates the second-best result.
@@
L2P [72] $3.94\pm 0.16$ $3.55\pm 0.20$ $5.01\pm 0.40$ $4.41\pm 0.43$ $10.76\pm 0.45$ $7.88\pm 0.17$
DualPrompt [71] $3.32\pm 0.16$ $2.78\pm 0.25$ $4.61\pm 0.07$ $3.70\pm 0.18$ $7.30\pm 0.18$ $5.16\pm 0.34$
CODA-P [60] $8.89\pm 0.65$ $7.65\pm 0.98$ $7.94\pm 0.08$ $6.72\pm 0.79$ $8.23\pm 0.86$ $6.95\pm 0.70$
LGCL [35] $3.04\pm 0.36$ $2.50\pm 0.38$ $4.75\pm 0.33$ $3.38\pm 0.58$ $7.35\pm 0.65$ $5.05\pm 0.32$
HiDe-Prompt [66] $3.15\pm 0.46$ $2.64\pm 0.16$ $\mathbf{2.29\pm 0.27}$ $\underline{2.33\pm 0.17}$ $\mathbf{2.23\pm 0.38}$
PGP [54] $3.36\pm 0.23$ $2.85\pm 0.25$ $4.53\pm 0.40$ $3.63\pm 0.35$ $7.17\pm 0.21$ $5.09\pm 0.25$
EvoPrompt [39] $\mathbf{1.79\pm 0.31}$ $\underline{1.41\pm 0.32}$ $4.22\pm 0.42$ $3.59\pm 0.52$ $6.72\pm 0.90$ $5.67\pm 0.26$
CPrompt [22] $6.00\pm 0.00$ $5.49\pm 0.00$ $6.10\pm 0.75$ $5.60\pm 1.35$ $5.98\pm 0.24$ $5.54\pm 0.48$
ConvPrompt [56] $3.42\pm 0.05$ $2.36\pm 0.16$ $4.17\pm 0.04$ $3.11\pm 0.17$ $\mathbf{4.87\pm 0.57}$ $3.57\pm 0.25$

### v3

@@ 4 Results
Comparison with PEFT CL approaches: As PEARL entails dynamic allocation of LoRA components, we compare PEARL with state-of-the-art PEFT CL methods across 5, 10, and 20 tasks in ImageNet-R [hendrycks2021many], in Table 2. The experiment is conducted using the ViT-B/16 backbone [dosovitskiy2020image] pre-trained on ImageNet-21k. As can be seen, PEARL (ViT) shows a significant improvement in both incremental and final accuracy, with absolute gains ranging from 9% to 14%. The performance improvement is more pronounced in 5 task sequence and slightly decreases thereafter as we move towards 10 and then to 20 task sequences. Among considered baselines, InfoLoRA is the closest work in terms of design and performance, but falls short of PEARL (ViT)’s performance by a huge margin across all scenarios. InfLoRA injects a small number of parameters to re-parameterize the pre-trained weights and shows that finetuning these injected parameters is equivalent to finetuning the pre-trained weights within a subspace. Since the reduction matrix $B$ infLoRA is handcrafted rather than being learnt from the training data, InfLoRA suffers a performance degradation compared to PEARL (ViT). Additional comparisons between our method and adapter-based as well as prompt-based approaches can be found in Table 14 and Table 14.
@@ A.2 Justification for considered baselines
To facilitate extensive evaluation of PEARL, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE); (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM); (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE); and (iv) PEFT approaches for continual learning (e.g., L2P, C-LoRA, InfLoRA). Firstly, inspired by the GWT, exemplar-based baselines entail a shared network for information reuse and task-specific attention modules that re-project the representation from the common representation space to the global workspace. Likewise, the PEARL framework includes a shared reference task network for information reuse and task-specific, layer-specific low-rank adapters for orienting towards the current task. Secondly, PEARL is a rehearsal-free framework; it does not store previous task samples either directly or indirectly through generative networks. Therefore, it aligns well with the NECIL benchmark and associated approaches. Thirdly, we note that low-rank adaptation introduces task-specific parameters that are trained on the corresponding task and kept frozen afterwards, akin to parameter-isolation approaches. As PEARL exhibits partial parameter isolation with a shared reference task network and low-rank adapters, we also provide a comparison with parameter isolation approaches. Finally, since PEARL readily utilizes LoRA [hu2021lora], a prominent PEFT technique, we offer an in-depth comparison with PEFT continual learning approaches. In addition, we also provide a discussion on related dynamic low-rank adaptation techniques beyond continual learning in Section A.3.
@@ A.3 Discussion on PEFT LoRA counterparts
PEFT CL approaches with static rank allocation: Among PEFT continual learning approaches, InfLoRA [liang2024inflora] and C-LoRA [DBLP:journals/corr/abs-2304-06027] are prominent techniques that involve low-rank adaptation of pre-trained networks, similar to PEARL. Each of these methods are tailored for CL scenarios and addresses unique challenges. InfLoRA aims to minimize the interference of new tasks on previously learned ones, thereby improving knowledge retention. However, its reliance on a handcrafted reduction matrix $B$, rather than a data-driven learned representation, limits its flexibility. C-LoRA, on the other hand, depends on a pre-trained model as its foundation, making it reliant on the availability of such models for its application. It applies LoRA to both the key and query attention mechanisms, ensuring efficient adaptation for new tasks. Additionally, C-LoRA incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. Despite its relative success, caution is warranted when training C-LoRA on large task sequences, as this approach suffers from scalability issues. In line with standard practices in continual learning benchmarks, we provide results for SeqLoRA, a sequential finetuning baseline without any safeguards against catastrophic forgetting. SeqLoRA experiences significant forgetting, underscoring the difficulty of maintaining performance across tasks in continual learning.
@@
PEFT approaches with dynamic rank allocation: Parameter efficient fine-tuning techniques such as LoRA distribute trainable parameters equally across all layers, which may reduce efficiency when layers vary in their contribution towards task adaptation. To address this, adaptive low-rank adaptation strategies have emerged that dynamically adjust parameter budgets to improve fine-tuning efficiency, similar to PEARL. AdaLoRA [zhang2023adaptive] adaptively allocates the parameter budget among weight matrices according to their importance score. The importance score is computed by parameterizing the incremental updates in the form of SV-decomposition and the singular values of unimportant updates are pruned. DyLoRA [valipour-etal-2023-dylora] proposed a dynamic low-rank adaptation in which LoRA blocks are trained for a range of ranks instead of a single rank by sorting the representation learned by the adapter module at different ranks during training. More recent methods, such as DoRA [mao-etal-2024-dora] and Dynamic Rank-Selective LoRA [lu2024adaptive], have emerged, with DoRA decomposing high-rank LoRA layers into structured single-rank components, while Dynamic Rank-Selective LoRA adaptively assigns ranks to LoRA modules based on their relevance to the current data in CL in vision language models. Although these approaches address a similar challenge as PEARL, these are not tailored for CL in custom, compact models.
@@ B.1 Comparison with fixed rank as a baseline
The PEFT CL approaches offer lightweight training schedules while effectively retaining the consolidated knowledge of pre-trained models. However, LoRA components in these methods are particularly sensitive to rank selection, which can lead to suboptimal resource allocation and diminished performance [valipour2022dylora]. Table 6 summarizes the average model size and performance metrics across various ranks for LoRA components. As the rank increases, performance improves at the expense of increased model size; however, beyond a certain rank, the performance gains do not really justify the additional increase in model size. Consequently, tuning the rank is essential to strike the right balance in this trade-off where model growth remains modest with optimum performance. To this end, PEARL employs a task-specific, layer-specific dynamic rank allocation enabling an optimal trade-off between model size and performance across different datasets and architectures.
@@ B.2 Comparison with convolutional parameter isolation approaches
Figure 2: Task-IL performance comparison with parameter isolation approaches on Seq-CIFAR100 (20T). We report the final Task-IL accuracies of each task after training on all tasks. Methods compared are NISPA [gurbuz2022nispa], PNN [rusu2016progressive], PAE [hung2019increasingly], CPG [hung2019compacting], PackNet [mallya2018packnet], and CLNP [golkar2019continual]
@@ B.5 Hyperparameter Tuning
Table 10: A benchmark comparison with CNN-based prior works on Class-IL ($A_{T}$% ) and Task-IL ($A_{T}$% ). The best results are highlighted in bold while the second best are underlined. The error terms indicate a standard deviation across 3 random runs. The rehearsal-based baselines employ a buffer size of 200.
@@
Class-IL ($A_{T}$% ) Task-IL ($A_{T}$% ) Class-IL ($A_{T}$% ) Task-IL ($A_{T}$% )
@@
PNNs [rusu2016progressive] NeurIPS 2016 - 74.01 $\pm$1.11 - 67.84 $\pm$0.29
PackNet [mallya2018packnet] CVPR 2018 - 72.39 $\pm$0.37 - 60.46 $\pm$1.22
NISPA [gurbuz2022nispa] NeurIPS 2022 - 65.36 $\pm$2.19 - 59.56 $\pm$0.32
SparCL-EWC [wang2022sparcl] NeurIPS 2022 - 59.53 $\pm$0.25 - 59.56 $\pm$0.32
@@
DER++ [buzzega2020dark] NeurIPS 2020 29.60 $\pm$1.14 62.49 $\pm$1.02 10.96 $\pm$1.17 40.87 $\pm$1.16
[…]
Co2L [cha2021co2l] ICCV 2021 31.90 $\pm$0.38 55.02 $\pm$0.36 13.88 $\pm$0.40 42.37 $\pm$0.74
CLS-ER [arani2022learning] ICLR 2022 43.80 $\pm$1.89 73.49 $\pm$1.04 23.47 $\pm$0.80 49.60 $\pm$0.72
OCDNet [li2022learning] IJCAI 2022 44.29 $\pm$0.49 73.53 $\pm$0.24 17.60 $\pm$0.97 56.19 $\pm$1.31
TAMiL [bhat2022task] ICLR 2022 41.43 $\pm$0.75 71.39 $\pm$0.17 20.46 $\pm$0.40 55.44 $\pm$0.52
IMEX-Reg [bhat_imex-reg_2024] TMLR 2024 48.54 $\pm$0.23 75.61 $\pm$0.73 24.15 $\pm$0.78 62.91 $\pm$0.54
TriRE [TriRE2023] NeurIPS 2023 43.91 $\pm$0.18 71.66 $\pm$ 0.44 20.14 $\pm$0.19 55.95 $\pm$0.78
AGILE [bhat2024mitigating] CoLLAs 2024 45.73 $\pm$0.15 74.37 $\pm$0.34 20.19 $\pm$1.65 53.47 $\pm$1.60
SARL [sarfrazsemantic] ICLR 2025 48.96 $\pm$ 0.53 78.91 $\pm$ 0.15 28.95 $\pm$ 1.13 71.56 $\pm$0.18
@@ B.7 Performance in longer task sequences
Figure 3: Evaluation of Class-IL ($A_{T}$% ) performance under longer task sequences in Seq-CIFAR100. In line with other parameter isolation approaches, there is a steady drop in performance as we move from 5 tasks to 10 tasks and eventually to 20 tasks.
@@ B.9 Dynamic rank allocation across layers in ViT
PEARL introduces a PEFT CL framework that incorporates dynamic rank selection for LoRA components during training. PEARL leverages reference task weights and adaptively determines the rank of task-specific, layer-specific LoRA components based on the proximity of the current task to the reference task weights in the parameter space. Figure 1 illustrates the dynamic rank allocation across layers for convolutional variants for the final task in Seq-CIFAR100 (5T). From left to right, Figure 1 presents a detailed overview of the layer-wise dynamic thresholds established based on task proximity, along with the corresponding selected rank and actual rank for our convolutional variants. Similarly, Figure 4 presents the dynamic rank allocation across layers for the final task in ImageNet-R (5T) for PEARL (ViT). In the ViT model, earlier layers are primarily responsible for extracting low-level image features. These features lay the foundation for subsequent high-level feature extraction and semantic understanding in later layers. As these features relatively less complex and can be re-used across tasks, PEARL puts less emphasis on them through lower rank allocation. On the other hand, intermediate layers handle more abstract and complex features, such as parts of objects and entire objects, necessitating higher rank allocation. Therefore, PEARL puts a major emphasis on intermediate layers with higher rank allocation. The top layers of the ViT model are designed to capture high-level semantic features and global contextual information. By later layers, the model has likely already captured the most significant semantic features relevant to the task, and additional layers contribute marginally smaller improvements to the final output quality. Therefore, PEARL allocates a lower threshold and a corresponding lower rank in later layers. In line with the aforementioned findings in zeng2024peeling, PEARL displays a bell-curve shaped dynamic rank allocation across layers thereby optimizing resource allocation and maximizing information reuse.
@@ B.13 Additional comparison with prompt-based methods
Table 15 provides an extensive comparison of prompt-based methods in the continual learning (CL) setting on ImageNet-R. Following HiDe-Prompt [wang2023hide], we evaluate our method using the final forgetting measure (FFM) and cumulative forgetting measure (CFM), defined as follows. Let $A_{i,t}$ denote the accuracy on the $i$-th task after learning the $t$-th task.
@@
Table 15: Following [wang2023hide], we compare additional prompt-based methods in the Class-IL setting on ImageNet-R using a ViT-B/16 backbone pre-trained on ImageNet-21K. Bold indicates the best result and underline indicates the second-best result.
@@
L2P [wang2022learning] $3.94\pm 0.16$ $3.55\pm 0.20$ $5.01\pm 0.40$ $4.41\pm 0.43$ $10.76\pm 0.45$ $7.88\pm 0.17$
DualPrompt [DBLP:conf/eccv/0002ZESZLRSPDP22] $3.32\pm 0.16$ $2.78\pm 0.25$ $4.61\pm 0.07$ $3.70\pm 0.18$ $7.30\pm 0.18$ $5.16\pm 0.34$
CODA-P [smith2023coda] $8.89\pm 0.65$ $7.65\pm 0.98$ $7.94\pm 0.08$ $6.72\pm 0.79$ $8.23\pm 0.86$ $6.95\pm 0.70$
LGCL [pub13491] $3.04\pm 0.36$ $2.50\pm 0.38$ $4.75\pm 0.33$ $3.38\pm 0.58$ $7.35\pm 0.65$ $5.05\pm 0.32$
[…]
PGP [qiao2024prompt] $3.36\pm 0.23$ $2.85\pm 0.25$ $4.53\pm 0.40$ $3.63\pm 0.35$ $7.17\pm 0.21$ $5.09\pm 0.25$
EvoPrompt [Kurniawan_Song_Ma_He_Gong_Qi_Wei_2024] $\mathbf{1.79\pm 0.31}$ $\underline{1.41\pm 0.32}$ $4.22\pm 0.42$ $3.59\pm 0.52$ $6.72\pm 0.90$ $5.67\pm 0.26$
CPrompt [gao2024consistent] $6.00\pm 0.00$ $5.49\pm 0.00$ $6.10\pm 0.75$ $5.60\pm 1.35$ $98\pm 0.24 5.5$$54\pm 0.48$
ConvPrompt [roy_2024_CVPR] $3.42\pm 0.05$ $2.36\pm 0.16$ $4.17\pm 0.04$ $3.11\pm 0.17$ $\mathbf{4.87\pm 0.57}$ $3.57\pm 0.25$